During one scene in the movie, X-Men: Days of Future Past, Magneto uses his magnetic powers to drag Mystique across the ground toward him, relying on a bullet which has been lodged in Mystique’s calf to pull her with a magnetic force, as shown below. Assume that Mystique has a mass of 61.2 kg and that the coefficient of kinetic friction between her and the ground is 0.280. With what force should Magneto pull her if the direction of his magnetic force is 30 degrees above the horizontal, and he wants to pull her toward him at a constant speed of 4 m/s? Give your answer in Newtons.
Answer #1

here,

mass of Mystique , m = 61.2 kg

the coefficient of friction , uk = 0.28

theta = 30 degree

let the force applied be F

the normal force , N = m * g - F * sin(theta)

for horizontal direction

for constant speed

the acceleration is zero

so , the net force is also zero

F * cos(theta) - uk * N = 0

F * cos(30) - 0.29 * ( 61.2 * 9.81 - F * sin(30)) = 0

solving for F

F = 172.21 N

the magnitude of force is 172.21 N